The Carma S. Diego Travellers Corp. has 10 million shares of stock outstanding at a current market price of $10. It is considering a new share offering that will net it $9 a share on 1 million shares. Earnings this year are expected to be $15 million. (Do not round intermediate calculations. Round the final answers to 2 decimal places.)

a. What is the immediate dilution potential for this new share issue? Dilution $

b-1. Assume Carma S. Diego Travellers Corp. can earn 10 percent on the proceeds of the share issue and these can be realized with this year’s results. Calculate earnings per share. Earnings per share $

b-2. Should the new issue be undertaken based on earnings per share? Yes No

c-1. If the 1 million additional shares can be issued to net $17 per share and the company can earn 10 percent on the proceeds, calculate earnings per share. Earnings per share $

c-2. Should the new issue be undertaken based on earnings per share? Yes
Answer #1

Requirement a.

Earnings per share before stock issue (original EPS) = 15,000,000 / 10,000,000 = $ 1.5

Earnings per share after stock issue ( Diluted EPS ) = 15,000,000 / 11,000,000 = $1.36364

Answer : Dilution per share = $ $ 1.5 - 1.36364 = $0.14

Requirement b -1 :

$9 x 1,000,000 shares x 0.10 = $ 900,000

Total earnings for the current year = 900,000 + 15,000,000 = 15,900,000

Earnings per share after additional income = 15,900,000 / 11,000,000 = $1.45

Requirement b-2 :

NO. New issue should not be undertaken because diluted EPS is less than the original EPS i.e., $1.45 is less than $1.5

Requirement C-1 :

$17 x 1,000,000 shares x 0.10 = $ 1,700,000

Total earnings for the current year = 1,700,000 + 15,000,000 = 16,700,000

Earnings per share after additional income = 16,700,000 / 11,000,000 = $1.52

Requirement C-2 :

Yes, the new issue should be undertaken as the new EPS is higher than original EPS i.e., $1.52 is higher than $1.5